OpenStudy (anonymous):

In the morning, Margret drove to a business appointment at 50 mph. Her average speed on the return trip in the afternoon was 40 mph. The return trip took (1/4) hours longer because of heavy traffic. How far did she travel to the appointment?

OpenStudy (anonymous):

50 km

OpenStudy (anonymous):

if she travelled x km, then \[x/40-x/50=1/4 \] or, \[x/200=1/4.\] hence you will get x as 50